math

A deck of cards contains 52 different cards. For each of the following situations, give the number of equally likely outcomes (a) I draw the top card. (b) I draw the top two cards. The order matters (c) I draw the top two cards. The order doesnt matter (d) I draw the top card. Without replacing the card, I reshuffle the remaining cards in the deck, and draw the top card. (The order matters.) (e) I draw the top card. Then I replace the card, reshuffle the deck, and draw the top card.
Answer #1

a) as there are 52 cards ; and selecting one card from 52 =52P1 =52 equally likely outcome

b)for order matters ; number of ways =52P2 =52*51 =2652

c) for order not matter ; number of ways =52C3 =22100

d) this is similar to problem 2 of having 2 cards from deck ; number of ways =52P2 =52*51 =2652

e) if order matters ;number of ways =52*52 =2704